OpenStudy (christina166):

The strength S of a rectangular beam varies directly as its width W and the square of its depth D, and inversely as its length L. Write an equation for this relationship.

OpenStudy (mathstudent55):

y varies directly as x: \(y = kx\) y varies inversely as z: \(y = \dfrac{k}{z} \) y varies directly as the square of w: \(y = kw^2\) Put all those together and find your answer.

OpenStudy (christina166):

what do you mean? im confused

OpenStudy (christina166):

@mathstudent55 how to i put this into an equation

OpenStudy (christina166):

@mathmate

OpenStudy (welshfella):

just bring them together Hint:- k only appears onc

OpenStudy (christina166):

wait how do i bring it together if k only appears once?

OpenStudy (christina166):

@welshfella how do i put it together if k appears once?

OpenStudy (anthonyym):

So what mathstudent55 wrote is an example. To apply it to your problem, you can see for example that y = kx is S = kW (for your problem).

OpenStudy (christina166):

oh okay

OpenStudy (christina166):

so the equation for the relationship will be y=kx/z

OpenStudy (christina166):

all i need to do is find the equation

OpenStudy (anthonyym):

So using mathstudent55's example can you figure out the others: The strength S of a rectangular beam varies directly the square of its depth D, and The strength S of a rectangular beam varies inversely as its length L.

OpenStudy (christina166):

so the strength S varies both directly to depth D and length L, i knew that but im confused on how the equation is going to be put together

OpenStudy (welshfella):

k only appears once in the equation also use w for width, D for depth and L for length

OpenStudy (anthonyym):

Yes, so what you should do is put them as an equation to make it easier (not word form). For example, The strength S of a rectangular beam varies directly as its width W becomes S = kW

OpenStudy (welshfella):

- sorry i was called away

OpenStudy (christina166):

so S(w)=kW?

OpenStudy (christina166):

im sorry i havent learned this technique yet and my teacher needs us to answer it

OpenStudy (mathmate):

@christina166 I find it a little difficult to follow when the expressions do not use the same notations as defined in the question. If I try to recap what was said, then, using the given notation in "The strength S of a rectangular beam varies directly as its width W and the square of its depth D, and inversely as its length L. Write an equation for this relationship" we have \(S\propto W\) \(S\propto D^2\) \(S\propto /L\) then \(S\propto WD^2/L\) where \(\propto\) reads : "proportional to" So it's your turn to remove the \(\propto\) and replace with a constant, as discussed earlier.

OpenStudy (christina166):

so the equation would end up being that but replaced with a constant

OpenStudy (christina166):

@mathmate so it would be \[S approxWD ^{2}/L\]

OpenStudy (mathmate):

It's \propto is you want to use LaTex. That would be ok, but probably your teacher wants to see a constant, usually k, as @mathstudent55 did in the first place.

OpenStudy (christina166):

so it would then be SkWD^2/L

OpenStudy (christina166):

THANK YOU !!!

OpenStudy (mathmate):

Would be perfect if there is an equal sign after "S".

OpenStudy (mathmate):

You're welcome!

OpenStudy (christina166):

yes S=kWD^2/L

OpenStudy (christina166):

OMG THANKS!!!!!!!!!!!!

OpenStudy (mathmate):

Yep, like that!

OpenStudy (mathstudent55):

If it varies directly as a variable, write the variable in the numerator. If it varies directly as the square of the variable, write the square of the variable in the numerator. If it varies inversely as a variable, write the variable in the denominator. Then include a constant of proportionality, such as k. \(\large S = \dfrac{kWD^2}{L} \)